# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v1

@@ 10 Conclusion
We presented a scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. Our results reveal that 58.7% of analyzed GPTs exhibit potential policy violations, highlighting significant gaps in the current publication and review process of the GPT store. Notably, compliance issues appear to stem primarily from the base models, such as GPT-4 and GPT-4o, rather than from user customizations. This indicates that foundational models still require improvements to better align with OpenAI’s usage policies while maintaining their versatility and utility.
Our framework, validated with an F1 score of 0.975, provides a robust tool for large-scale, systematic policy compliance evaluation. Given its scalability and effectiveness, we propose its integration into OpenAI’s review process at the time of publication of Custom GPTs in the GPT store. This addition would enhance the oversight of model compliance, reducing the prevalence of unsafe or non-compliant deployments.
As future work, we aim to extend the validation of our framework to include experts in a broader range of usage policies and application domains, such as healthcare, finance, and legal contexts. While deceptive prompts were excluded from the large-scale evaluations due to challenges in annotator agreement, they hold significant potential to enhance compliance detection in complex scenarios. Future efforts will focus on refining these prompts and achieving consistent annotations, enabling their integration into the framework alongside direct prompts. We also plan to explore more sophisticated adversarial testing methods, such as multi-turn interactions and prompt injection strategies, to uncover deeper and broader compliance issues. Together, these advancements will further refine the framework’s utility and contribute to the broader goal of ensuring the safety of large language models and personalized chatbots.
Open Science
In line with principles of transparency, reproducibility, and fostering collaboration within the research community, we plan to release the following resources to a public repository [^2] The repository will be made public upon paper acceptance.:
@@
Appendix: Extended OpenAI Policies
@@
Policy for Romantic GPTs
@@
Policy for Cybersecurity GPTs
@@
Policy for Academic GPTs

### v2

@@ 11 Conclusion
We presented a scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. Our results reveal that 58.7% of analyzed GPTs exhibit potential policy violations, highlighting significant gaps in the current publication and review process of the GPT store. Notably, compliance issues appear to stem primarily from the base models, such as GPT-4 and GPT-4o, rather than from user customizations. This indicates that foundational models still require improvements to better align with OpenAI’s usage policies while maintaining their utility.
Our framework, achieving an F1 score of 0.975, is an effective tool for large-scale, systematic policy compliance evaluation. Given its scalability and efficacy, we propose its integration into OpenAI’s review process for Custom GPTs prior to their publication in the GPT Store. This addition would enhance the oversight of model compliance, reducing the prevalence of unsafe or non-compliant deployments.
As future work, we plan to extend the validation of our framework to include experts across additional policy areas and sensitive domains such as healthcare, finance, and law. Although deceptive prompts were excluded from the large-scale evaluation due to annotation challenges, they remain promising for detecting complex compliance issues. We aim to refine these prompts and improve annotation consistency to incorporate them into future assessments. Additionally, we will explore advanced adversarial strategies, including multi-turn interactions and prompt injection, to uncover broader risks. These efforts will enhance the framework’s utility and contribute to safer LLM-based systems.
Appendix A Open Science
In line with principles of transparency, reproducibility, and fostering collaboration within the research community, we will publicly release the following resources upon acceptance of the paper:
@@ Appendix A Open Science
Appendix B Ethics Considerations
This study raises important ethical considerations stemming from the automation of interactions with OpenAI’s GPT Store. While the proposed framework provides valuable insights into compliance issues in Custom GPTs, its implementation involves challenges. This section outlines the ethical implications of these challenges, discusses their alignment with the principles of the Menlo Report, and describes the mitigation strategies adopted to address them.
Automation and Terms of Service
The GPT Collector & Interactor module retrieves metadata, generates prompts, and collects responses from Custom GPTs via the GPT store’s graphical user interface (GUI). While necessary for the large-scale evaluation conducted in this study, such automation contravenes OpenAI’s terms of service (ToS). This issue touches on the Respect for Law and Public Interest principle outlined in the Menlo Report, which emphasizes compliance with established rules and policies to uphold trust and sustainability in digital ecosystems.
The decision to automate interactions via the GUI was driven by the absence of an official API or other means to programmatically access and evaluate Custom GPTs. While this approach contradicts OpenAI’s terms of service, it enables contributions to the research community. Specifically, it facilitated the large-scale evaluation of compliance in Custom GPTs, highlighted significant safety risks for users and society, and exposed governance weaknesses in the GPT Store that warrant further scrutiny.
However, releasing the GPT Collector & Interactor module as part of the framework poses significant risks. Such a release could enable misuse by malicious actors, including large-scale scraping of GPT metadata, spam-like automation, and circumventing OpenAI’s API infrastructure, resulting in financial harm and operational strain on the platform. To mitigate these risks, we have chosen to exclude this module from the publicly available framework code. Instead, access will be granted to researchers upon request, contingent on verification of their research position and objectives. This decision reflects our commitment to responsible research dissemination, ensuring that the benefits of our work can be realized without creating undue harm to OpenAI’s infrastructure.
Live System Interactions
The framework’s reliance on querying live systems introduces additional ethical considerations. Directly interacting with OpenAI’s GPT Store to retrieve metadata and conduct compliance evaluations imposes operational costs and risks disrupting platform services. These interactions could inadvertently affect other users sharing the same infrastructure, raising concerns under the Beneficence principle, which emphasizes minimizing harm to stakeholders while maximizing societal benefits. To mitigate these risks, the queries were carefully limited in scope and frequency to the rates imposed by OpenAI, ensuring minimal impact on the platform.
Responsible Disclosure to OpenAI
As part of our commitment to ethical research practices, we proactively disclosed our methodology and findings to OpenAI, as disclosed in §7. Their response acknowledged the value of our automated evaluation approach for identifying compliance issues, signaling recognition of the implications of our study.
This disclosure reflects our alignment with the Respect for Law and Public Interest principle of the Menlo Report. By engaging transparently with OpenAI, we aimed to provide them with early insights into our findings, enabling them to address systemic issues or vulnerabilities in the GPT Store before potential misuse, thus contributing to the ethical responsibility of reducing risks to the platform and its users. We hope this engagement strengthens platform governance and compliance mechanisms.
Appendix C Operationalized OpenAI Policies
@@
C.1 Policy for Romantic GPTs
@@
C.2 Policy for Cybersecurity GPTs
@@
C.3 Policy for Academic GPTs